OpenStudy (cutestgirlaround):

Please help!!! solve for w? −(1/4+8)+8=3 A. W= -12 B. W= - 3/4 C. W= 13/4 D. W= 52

OpenStudy (cutestgirlaround):

−(1/4w+8)+8=3 sorry for the confusion

Awolflover1 (awolflover1):

\[( \frac{ -1 }{ 4 } w)+(−8+8)=3\] \[\frac{ -1 }{ 4 } w=3\]

Awolflover1 (awolflover1):

Multiply both sides by -4/1 \[( \frac{- 4 }{ 1 } )\times( \frac{ -1 }{ 4} w)=(\frac{ -4 }{ 1 } )\times3)\]

OpenStudy (cutestgirlaround):

-12???

Awolflover1 (awolflover1):

yes :)

OpenStudy (cutestgirlaround):

Thanks!!!

OpenStudy (cutestgirlaround):

I need more help so can I post them so you can try and help

Awolflover1 (awolflover1):

alright =)